# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

### v1

@@ 7       Set $a_{t+1}=\nabla\mathcal{R}_{t}^{-1}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t})$
$D_{\mathcal{R}^{*}_{t}}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t},\nabla\mathcal{R}_{t}(a_{t}))\leq 4d^{2}\eta^{2}.$
Lemma C.15 (Lemma 4 in Hazan and Levy (2014)).
@@
If we let $C_{0}=(2\nu+4L_{0}\beta+4B_{2}L^{2}+LL_{0}\beta)/\lambda$, we have
$\text{Reg}^{\text{DB}}_{T}$ $\leq\frac{2\nu\log(T)}{\eta_{\rho}}+8d^{2}\eta_{\rho}T+2\mathbb{E}\left\{\sum^{T}_{t=1}b_{t}^{\top}(a_{t}-a^{*}_{T})\right\}+\frac{4L_{0}\beta+4B_{2}L^{2}+LL_{0}\beta}{\lambda\eta_{\rho}}\log T+2LL_{0}R$
[…]
@@
C.3.4 Proof for Lemma C.10: Observation Error
Proof  (Proof for Lemma C.10) In the following, we will analyze the expected cumulative sum of the bias $b_{t}$ on the total regret, which is $\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$. By Cauchy-Schwarz inequality, $b^{\top}_{t}(a_{t}-a^{*}_{T})\leq\|b_{t}\|_{2}\|(a_{t}-a^{*}_{T})\|_{2}$. We can bound the $\|b_{t}\|_{2}$ by the following.
@@
We get the first equality by Lemma C.8. We get the inequality by using the fact $\|u_{t}\|_{2}=1$. If we let $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}$, then we have
@@
$\leq\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right).$
We get the first equality by using the definition of $P_{t}(a_{t})$ and $\hat{P}_{t}(a_{t})$. We get the first inequality by using the Lipschitz property of $\sigma$ (see Assumption C.3) Therefore, we have
$\|b_{t}\|_{2}\leq d\sqrt{\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))}\min\left\{2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right\}.$
@@
$\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))$ $=\lambda_{\max}(\nabla^{2}\mathcal{R}(a_{t}))+\lambda\eta_{\rho}t+2\mu\leq\lambda^{*}_{R}+2\mu+\lambda\eta_{\rho}t.$
@@
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right))\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda^{*}_{R}+2\mu+\lambda\eta_{\rho}t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right))\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda^{*}_{R}+2\mu}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}+\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda\eta_{\rho}t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$ (23)
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+2d\sqrt{\lambda\eta_{\rho}T}+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}\|_{2}\right\}$ (24)
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+2d\sqrt{\lambda\eta_{\rho}T}+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\min\left(2R,\sqrt{\frac{2}{\alpha}\left(f(a_{t})-f(a^{*})\right)}\right)\right\}.$
We get Equation (23) by using the fact that $\|a_{t}-a^{*}_{T}\|_{2}\leq 2R$, $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq C$. We get Equation (24) the definition of $a^{*}_{T}$, and the fact $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$. Using the $\alpha$-strong convexity of $f$ (see Assumption C.2), we have $\|a_{t}-a_{*}\|_{2}\leq\min\left(2R,\sqrt{\frac{2}{\alpha}\left(f(a_{t})-f(a^{*})\right)}\right)$ we get the last inequality, which completes the proof.
C.3.5 Proof for Lemma C.11
Proof  (Proof for Lemma C.11). Using Abel’s Summation Equation (Lemma C.16), we have
@@
Since we have $\sum^{T}_{t=1}n^{2}_{t}\leq C^{\prime}T^{\frac{1}{2}+\alpha}$ and $0\leq n_{t}\leq K,\forall t$, we have $\sum^{T}_{t=1}n_{t}\leq\sqrt{C^{\prime}}T^{\frac{3}{4}+\frac{\alpha}{2}},\forall t$. Moreover, for $t\geq 1$, $t^{-1+\rho}-(t+1)^{-1+\rho}\leq t^{\rho-\frac{1}{2}}$ holds for all $t\geq 1$. This is because
@@
$\leq\sqrt{C^{\prime}}c_{k}T^{\frac{3}{4}+\frac{\alpha}{2}}T^{-1+\rho}+\sqrt{C^{\prime}}c_{k}\sum^{T-1}_{t=1}\left(\sum^{t}_{i=1}n_{i}\right)t^{-\frac{1}{2}+\rho}$
$\leq\sqrt{C^{\prime}}c_{k}T^{\frac{3}{4}+\frac{\alpha}{2}}T^{-1+\rho}+\sqrt{C^{\prime}}c_{k}\sum^{T-1}_{t=1}t^{\frac{3}{4}+\frac{\alpha}{2}}t^{-\frac{1}{2}+\rho}$
@@
C.4 Proof for Proposition 2: Efficiency-Robustness Tradeoff in NC-SMD
Theorem (Full Statement of Proposition 2).
Let the number of rounds $T$ be fixed and $T>\sqrt{2LL_{0}R}$. Under Assumptions C.2, C.3, and C.4, by choosing $\lambda\leq l_{0}\alpha/2$, $\mu\geq(L^{3}_{0}L_{2}/\lambda)^{2}$, and setting the learning parameter $\eta_{\rho}=\frac{\sqrt{\log T}}{2d}T^{-\alpha},\alpha\in[0.5,1]$, regret occurred by Algorithm 3 satisfies
$\text{Reg}_{T}\leq\tilde{O}\left(dT^{\alpha}+\sqrt{d}T^{\frac{1}{2}(1-\alpha)}C+dC\right)$
for any corruption with budget $C$.
Proof  (Proof for Proposition 2). From Lemma C.10, we know
@@ C.4 Proof for Proposition 2: Efficiency-Robustness Tradeoff in NC-SMD
$\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$ $\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+2Rd\sqrt{\lambda\eta_{\rho}}L_{\sigma}\sum^{T}_{t=1}\sqrt{t}|c_{t}(a_{t},a^{\prime}_{t})|$
$\leq 2RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+2Rd\sqrt{\lambda\eta_{\rho}T}L_{\sigma}C.$
We get the second inequality by using the fact that the diameter of the action space of $R$. We get the last inequality by using Abel Summation Equation (see Lemma C.16). Therefore, by Lemma C.9, we have
$\text{Reg}_{T}\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2LL_{0}R+4RdL_{\sigma}\sqrt{\lambda^{*}_{R}+2\mu}C+4Rd\sqrt{\lambda\eta_{\rho}T}L_{\sigma}C.$

### v2

@@ B Proof for Theorem 2
$D_{\mathcal{R}^{*}_{t}}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t},\nabla\mathcal{R}_{t}(a_{t}))\leq 4d^{2}\eta^{2}_{\rho}.$
Lemma 19 (Lemma 4 in Hazan and Levy (2014)).
@@
If we let $C_{0}:=(2\nu+4L^{\sigma}_{1}\kappa+4R^{\sigma}_{2}(L^{\mu})^{2}+L^{\mu}L^{\sigma}_{1}\kappa)/\lambda$, we have
$\text{Reg}_{T}$ $\leq\frac{2\nu\log(T)}{\eta_{\rho}}+8d^{2}\eta_{\rho}T+2\mathbb{E}\left\{\sum^{T}_{t=1}b_{t}^{\top}(a_{t}-a^{*}_{T})\right\}+\frac{4L^{\sigma}_{1}\kappa+4R^{\sigma}_{2}(L^{\mu})^{2}+L^{\mu}L^{\sigma}_{1}\kappa}{\lambda\eta_{\rho}}\log T+2L^{\mu}L^{\sigma}_{1}R$
$=\underbrace{\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2L^{\mu}L^{\sigma}_{1}R}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Regret\hskip 2.84526ptof\hskip 2.84526ptDecision}}+\underbrace{2\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Feedback\hskip 2.84526ptError}},$
@@
Step 3: upper bounding the regret from feedback error
Bounding the regret of decision term in (2) is standard. Thus, this step develops a novel upper bound for bounding the regret from feedback error, as shown below.
Lemma (Lemma 4 restated).
The feedback error term in (2) can be bounded as follows:
$\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$ $\leq C_{1}d\sqrt{\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}|c_{t}(a_{t},a^{\prime}_{t})|\sqrt{\mu(a^{*})-\mu(a_{t})}\right\}+C_{2}dT^{\rho}+2d\sqrt{\lambda}\sqrt{\eta_{\rho}T},$
where $C_{1}:=\sqrt{\frac{2\lambda(L^{\sigma})^{2}}{\alpha}}$, and $C_{2}:=2L^{\sigma}R\sqrt{\sup_{a\in\mathcal{A}}\lambda_{\max}\left(\nabla^{2}\mathcal{R}(a)\right)+2\phi}$.
Proof  In the following, we will analyze the expected cumulative sum of the bias $b_{t}$ on the total regret, which is $\mathbb{E}\left\{\sum^{T}_{t=1}b^{\top}_{t}(a_{t}-a^{*}_{T})\right\}$. By Cauchy-Schwartz inequality, $b^{\top}_{t}(a_{t}-a^{*}_{T})\leq\|b_{t}\|_{2}\|(a_{t}-a^{*}_{T})\|_{2}$. We can bound the $\|b_{t}\|_{2}$ by the following.
@@
We get the first equality by Lemma 2. We get the inequality by using the fact $\|u_{t}\|_{2}=1$. If we let $a^{\prime}_{t}=a_{t}+\nabla^{2}\mathcal{R}_{t}(a_{t})^{-\frac{1}{2}}u_{t}$, then we have
@@
$\leq\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right).$
We get the first equality by using the definition of $P_{t}(a_{t})$ and $\hat{P}_{t}(a_{t})$. We get the first inequality by using the Lipschitz property of $\sigma$. Therefore, we have
$\|b_{t}\|_{2}\leq d\sqrt{\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))}\min\left\{2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right\}.$
@@
$\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))$ $=\lambda_{\max}(\nabla^{2}\mathcal{R}(a_{t}))+\lambda\eta_{\rho}t+2\phi\leq\lambda^{*}_{R}+2\phi+\lambda\eta_{\rho}t.$
@@
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda_{\max}(\nabla^{2}\mathcal{R}_{t}(a_{t}))}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right))\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda^{*}_{R}+2\phi+\lambda\eta_{\rho}t}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right))\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda^{*}_{R}+2\phi}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}+\mathbb{E}\left\{\sum^{T}_{t=1}(d\sqrt{\lambda\eta_{\rho}t}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$
$\leq 2RdL^{\sigma}\sqrt{\lambda^{*}_{R}+2\phi}T^{\rho}+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}_{T}\|_{2}\right\}$ (16)
$\leq 2RdL^{\sigma}\sqrt{\lambda^{*}_{R}+2\phi}T^{\rho}+2d\sqrt{\lambda\eta_{\rho}T}+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\|a_{t}-a^{*}\|_{2}\right\}$ (17)
$\leq 2RdL^{\sigma}\sqrt{\lambda^{*}_{R}+2\phi}T^{\rho}+2d\sqrt{\lambda\eta_{\rho}T}$
$+d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)\min\left(2R,\sqrt{\frac{2}{\alpha}\left(\mu(a^{*})-\mu(a_{t})\right)}\right)\right\}.$ (18)
We get (16) by using the fact that $\|a_{t}-a^{*}_{T}\|_{2}\leq 2R$, $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq T^{\rho}$. We get (17) the definition of $a^{*}_{T}$, and the fact $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$. Using the $\alpha$-strong convexity of $\mu$, we have $\|a_{t}-a_{*}\|_{2}\leq\min\left(2R,\sqrt{\frac{2}{\alpha}\left(\mu(a^{*})-\mu(a_{t})\right)}\right)$ we get the last inequality, which completes the proof.
Step 4: The iterative regret refinement analysis and resultant optimal learning rate
Lemma (Lemma 5 restated).
If $A=\mathbb{E}\left(\sum^{T}_{t=1}[\mu(a^{*})-\mu(a_{t})]\right)\leq\mathcal{O}(T^{\frac{1}{2}+\psi})$ for some $\psi\in[0,\frac{1}{2})$, then we must have $B=\mathbb{E}\left(\sum^{T}_{t=1}t^{\rho-\frac{1}{2}}\sqrt{\mu(a^{*})-\mu(a_{t})}\right)\leq\mathcal{O}(T^{-\frac{1}{4}+\frac{\psi}{2}+\rho})$.
Proof  To prove Lemma 5, it is equivalent to prove the following Lemma 20. If Lemma 20 holds, let $n_{t}=\sqrt{\mathbb{E}(\mu(a^{*})-\mu(a_{t}))}$, $m_{t}=\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)$, proves Lemma 5.
Lemma 20.
Proof
Using Abel’s equality (Lemma 21), we have
@@
Since we have $\sum^{T}_{t=1}n^{2}_{t}\leq C^{\prime}T^{\frac{1}{2}+\alpha}$ and $0\leq n_{t}\leq K,\forall t$, we have $\sum^{T}_{t=1}n_{t}\leq\sqrt{C^{\prime}}T^{\frac{3}{4}+\frac{\alpha}{2}},\forall t$. Moreover, for $t\geq 1$, $t^{-1+\rho}-(t+1)^{-1+\rho}\leq t^{\rho-2}$ holds for all $t\geq 1$. This is because
@@
$\leq\sqrt{C^{\prime}}c_{k}T^{\frac{3}{4}+\frac{\alpha}{2}}T^{-1+\rho}+\sqrt{C^{\prime}}c_{k}\sum^{T-1}_{t=1}\left(\sum^{t}_{i=1}n_{i}\right)t^{-2+\rho}$
$\leq\sqrt{C^{\prime}}c_{k}T^{\frac{3}{4}+\frac{\alpha}{2}}T^{-1+\rho}+\sqrt{C^{\prime}}c_{k}\sum^{T-1}_{t=1}t^{\frac{3}{4}+\frac{\alpha}{2}}t^{-2+\rho}$
@@
After establishing Lemma 5, we can use it to prove the induction claim (Lemma 6). The challenge in this proof lies in identifying the constant $K$, which must hold across all iterations of the analysis to ensure the bound remains stable and does not diverge. This requires careful refinement of the analysis and calculations.
Lemma (Lemma 6 restated).
Consider $T>\sqrt{2L^{\mu}L^{\sigma}_{1}R}$ and $\eta_{\rho}=\frac{\sqrt{\log T}}{dT^{\max\{1/2,\rho\}}}$. If $\text{Reg}_{T}\leq 144dK\sqrt{\log T}T^{\rho+\frac{\frac{3}{2}-\rho}{2^{k}}}$ for some integer $k$, then we must also have $\text{Reg}_{T}\leq 144dK\sqrt{\log T}T^{\rho+\frac{\frac{3}{2}-\rho}{2^{k+1}}}$. $K$ is a instance-dependent constant, where $K:=\max\left\{C_{0},\frac{C_{2}C_{\kappa}}{\sqrt{\log T}},(L^{\sigma}T^{\rho}_{\kappa})^{2},2\sqrt{\lambda}RC_{\kappa}L^{\sigma}\right\}$, $C_{0}$ and $C_{2}$ defined in Lemma 3 and 4 respectively,
Proof  In the following, we denote
$m_{t}:=\min\left(2,L^{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right),n_{t}:=\min\left(2R,\sqrt{\frac{2}{\alpha}\left(\mu(a^{*})-\mu(a_{t})\right)}\right)$
Continue from (18), we have the following.
The Base Case: When $k=1$, let $C_{3}=C_{2}dT^{\rho}+2d\sqrt{\lambda\eta_{\rho}T}$, we have
@@
$\text{Reg}_{T}$ $\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2L^{\mu}L^{\sigma}_{1}R+2C_{3}+2d\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left(\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right).$ (19)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2L^{\mu}L^{\sigma}_{1}R+2C_{3}+4dR\sqrt{\lambda\eta_{\rho}}\mathbb{E}\left\{\sum^{T}_{t=1}\sqrt{t}m_{t}\right\}$ (20)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2L^{\mu}L^{\sigma}_{1}R+2C_{3}+4dR\sqrt{\lambda\eta_{\rho}T}\left(\sum^{T}_{t=1}m_{t}\right)$ (21)
$\leq\frac{C_{0}}{\eta_{\rho}}\log(T)+8d^{2}\eta_{\rho}T+2L^{\mu}L^{\sigma}_{1}R+2C_{3}+4dL^{\sigma}R\sqrt{\lambda\eta_{\rho}}C_{\kappa}T^{\rho}.$ (22)
We get (19) according to Lemma 3 and 4. We get (20) because of $\mathbb{E}\left(n_{t}\right)\leq 2R$. We get (21) by Abel’s Summation Equation (see Lemma 21). We get (22) because of corruption budget. Choosing the learning rate $\eta_{\rho}$ according to Theorem 2, we obtain $\text{Reg}_{T}\leq 144\sqrt{\log T}KdT^{\frac{\rho+1}{2}}$. This confirms the validity of the claim when $k=1$.
The Induction Argument: Let’s assume that the claim holds true for a general step $k$. In the following, we will demonstrate that the claim also holds true for step $k+1$. Because of Equation (14) and the induction claim, we obtain $\text{Reg}^{\text{FO}}_{T}\leq\frac{144}{l^{\sigma}_{1}}dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k}}}.$ This suggests that
$\sum^{T}_{t=1}\mathbb{E}\left(\mu(a^{*})-\mu(a_{t})\right)\leq\frac{144}{l^{\sigma}_{1}}dK\sqrt{\log T}T^{\rho+\frac{3/2-\rho}{2^{k}}}.$
By Jensen’s inequality, we have
$\mathbb{E}(n_{t})=\min\left\{2,\sqrt{\frac{2}{\alpha}}\mathbb{E}\left(\sqrt{\mu(a^{*})-\mu(a_{t})}\right)\right\}\leq\sqrt{\frac{2}{\alpha}}\sqrt{\mathbb{E}\left(\mu(a^{*})-\mu(a_{t})\right)}.$
Therefore we have
$\mathbb{E}\left(\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\right)\leq\sqrt{\frac{2}{\alpha}}\sum^{T}_{t=1}\sqrt{t}m_{t}\sqrt{\mathbb{E}\left(\mu(a^{*})-\mu(a_{t})\right)}.$
By the definition of $\rho$-Imperfect Human Feedback, we have $|c_{t}(a_{t},a^{\prime}_{t})|$ is upper bounded by $C_{k}t^{-1+\rho}$.
Therefore, using Lemma 5 we can upper bound $\sqrt{\lambda}\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}$ by
$\sqrt{\lambda}\sum^{T}_{t=1}\sqrt{t}m_{t}n_{t}\leq 60L^{\sigma}\sqrt{\frac{2\lambda}{\alpha l^{\sigma}_{1}}}(\log T)^{0.25}C_{\kappa}\sqrt{dK}T^{\frac{3}{2}\rho+\frac{3/2-\rho}{2^{k+1}}}$